OpenStudy (anonymous):

using implicit differentiation, find the derivative of x^2y + 3xy^3 - x = 3?

OpenStudy (amistre64):

just use normal differentiation :) theres no difference

OpenStudy (amistre64):

\[x^2y + 3xy^3 - x = 3\] \[\frac{d(x^2y)}{dx} + \frac{d(3xy^3)}{dx} - \frac{d(x)}{dx} = \frac{d(3)}{dx}\]

OpenStudy (amistre64):

the first term uses the product rule; the second term uses it as well the third is and last are simple enough

OpenStudy (anonymous):

Thanks for the help. We are required to do it by implicit differentiation is the reason I ask. :)

OpenStudy (amistre64):

implicit simply means that the y variable cannot be factored out and solved for; the rules for their derivates is the exact same procedures

OpenStudy (amistre64):

\[Dx(x^2y) = Dx(x^2)y + x^2Dx(y)\]

OpenStudy (amistre64):

\[Dx(3xy^3)=Dx(3x)y^3+3xDx(y^3)\]

OpenStudy (anonymous):

Kind of making sense now.

OpenStudy (anonymous):

Thank you.

OpenStudy (amistre64):

youre welcome :) dont the the variable throw you; the rules for derivatives doesnt care what the variable is power rule is power rule; chain rule is chain rule; product rule is product rune .... etc